-3x+3y=-3 x+y+-11 what do i do to solve for x and y

Answer by josgarithmetic(39617)

This is not the way to show your equations:
-3x+3y=-3 x+y+-11


THIS is how to show them:
-3x+3y=-3 
x+y=-11